# Montréal, QC Canada

### v1

@@ Abstract
Dynamic Mode Decomposition (DMD) is a numerical method that seeks to fit timeseries data to a linear dynamical system. In doing so, DMD decomposes dynamic data into spatially coherent modes that evolve in time according to exponential growth/decay or with a fixed frequency of oscillation. A prolific application of DMD has been to video, where one interprets the high-dimensional pixel space evolving through time as the video plays. In this work, we propose a simple and interpretable motion detection algorithm for streaming video data rooted in DMD. Our method leverages the fact that there exists a correspondence between the evolution of important video features, such as foreground motion, and the eigenvalues of the matrix which results from applying DMD to segments of video. We apply the method to a database of test videos which emulate security footage under varying realistic conditions. Effectiveness is analyzed using receiver operating characteristic curves, while we use cross-validation to optimize the threshold parameter that identifies movement.
@@
dynamic mode decomposition motion detection background subtraction computer vision
@@ 1 Introduction
Dynamic Mode Decomposition (DMD) is a method for forecasting and extracting complicated spatio-temporal behaviour from high-dimensional datasets. First developed by Schmid schmid2010dynamic , DMD seeks to identify a matrix $A$ so that advancing the data forward by one timestep is equivalent to multiplication by this matrix tu2014dmd ; DMDbook . In this way, DMD poses complex movement over space and through time as a linear discrete dynamical system via the matrix $A$. Since linear systems are entirely understood by their spectrum, the eigenvectors of $A$, called the DMD modes, can be seen to represent macroscopic patterns in the dataset that persist through time, either oscillating, growing, or decaying at a fixed rate determined by the associated eigenvalue. DMD has connections to the Koopman operator rowley2009spectral ; brunton2022modern and has proven to be a useful tool in identifying coherent structures in fluid mechanics schmid2011applications ; menon2020dynamic ; colbrook2023residual and neuroscience brunton2016extracting ; partamian2023analysis ; shiraishi2020neural , as well as to develop financial trading strategies mann2016dynamic , among numerous other applications; see DMDbook ; schmid2022dynamic and the references therein. Improvements on the method include building in known physics baddoo2023physics ; morandin2023port , accounting for irregular timesteps and noisy measurements lee2023optimized ; askham2018variable ; hemati2017biasing , and to incorporate the effect of control proctor2016dynamic ; garcia2022extended .
The utility of DMD has also made its way into computer vision. This includes employing DMD for edge detection in images bi2017dynamic , salient region detection sikha2021dynamic ; sikha2018salient ; sikha2020multi , and performing image classification rahul2019dynamic . Most important to the work herein, DMD has been applied to separate out the background and foreground in videos, allowing for object tracking krake2022efficient ; grosek2014dynamic ; kutz2015multi ; erichson2019compressed ; haq2020dynamic ; han2022quaternion . Roughly, this comes from the fact that eigenvectors of the DMD matrix $A$ with eigenvalues having modulus near 1 represent patterns in the (vectorized) video data that exhibit little change from frame to frame. Thus, these eigenvectors can be seen as representing the background of the video data, while the remaining eigenvectors fill in the foreground motion. Importantly, projecting the video onto the background/foreground DMD modes constitutes a separation of these distinct components of the video. The result is then two videos: one of a nearly static background and another of just the isolated motion, as exemplified in Figure 1.
@@
Figure 1: Screenshots of DMD being be used to separate out the background and foreground in video data. The video is taken from the Background Models Challenge database BMC .
@@
Figure 2: A visual overview of real-time motion detection and tacking with DMD. Sequential collections of video frames, or windows, are fed in and compressed to reduce computational complexity. Then, DMD is applied to identify spatio-temporal features within the windows. Spikes in the spectrum of the DMD matrices over the windows indicate movement in the video, thus detecting events as they occur in real-time.
@@
Our contribution in this paper is to present a method of real-time motion detection in streaming video data using DMD. To do this we apply DMD not to the entire video all at once, but to short windows as they are fed in. Precisely, we employ sliding window DMD dylewsky2019dynamic ; bramburger2020sparse to streaming video data to examine the dominant timescales in each segment of the video. These timescales are revealed by examining the modulus of the logarithm of the eigenvalues for the DMD matrix within each window. Sudden spikes in the DMD spectrum denote fast timescale movement in the foreground, allowing for motion detection. A visual overview of our method can be found in Figure 2, while a proper explanation of the method can be found in Section 3 below. Once movement has been detected, one can then use the DMD matrix to isolate the foreground motion by subtracting out the background. Computation time is accelerated by compressing the frames of the video erichson2019compressed , resulting in a computationally cheap and efficient method of real-time motion detection for streaming video data.
@@
Motion detection is a classic problem in computer vision with many complementary and competing methods. The simplest method identifies temporal differences cheung2005robust ; yu2009real , amounting to searching for significant change in the pixels from one frame to the next. This method is sensitive to abrupt changes in illumination, shadows, and repetitive motions, while also having the drawback that it cannot accurately extract the moving object boundaries. Alternatively, the work tsai2008motion transforms the frames into Fourier space where object movement can be shown to create a non-vertical structure in the 2D Fourier transform. While again being an effective method, it suffers from a high cost of computation. Computation cost can be brought down using the cosine transform oh2012real , for which graph cuts can be employed to isolate the movement in the video. Graph cuts come with their own well-documented limitations (see boykov2006graph and the references therein), including significant memory usage. Another drawback of coupling motion detection algorithms with graph cuts is the simple fact that identifying and isolating movement becomes a two-step task. A major advantage of our method herein is that DMD provides the necessary tools to both identify movement (through eigenvalues) and isolate it (through eigenvectors).
@@
The review singh2020object compares eight methods for real-time video surveillance of varying complexity. It is found that while simpler methods, such as examining temporal differences, are fast, they often fail to handle many of the complexities encountered in real-world video data. As expected, better performing methods come with higher computation time and a more complicated implementation. Our DMD-based method herein offers simplicity in its implementation, while also being backed up by dynamical systems theory to justify its performance. Fundamentally, DMD is a method of multiresolution analysis that accounts for spatial and temporal coherencies simultaneously. It identifies macroscopic patterns in the data and is more robust to scale overlap than related methods dylewsky2019dynamic , making it optimally suited for the task of motion detection.
@@
The paper is organized as follows. In Section 2 we provide an overview of DMD (§ 2.1, 2.2, 2.3) and its capabilities for background subtraction (§ 2.4), followed by a presentation of our own technique for motion detection in video (§ 3). In Section 4, we apply this method to a dataset of videos (§ 4.1) and discuss its performance (§ 4.2). The performance is then is evaluated through the lens of receiver operating characteristic curves (§ 4.3), and a strategy to optimize a key parameter involving $k$-fold cross-validation is presented (§ 4.4), along with concluding remarks in Section 5.
@@ 2 Background on Dynamic Mode Decomposition
In this section we provide an overview of all the relevant aspects of DMD for our method. We begin in § 2.1 with a presentation of DMD in its simplest form. We then turn to a review of DMD with compression in § 2.2, sliding window DMD in § 2.3 and background subtraction in § 2.4.
@@ 2.1 Dynamic Mode Decomposition
We begin with a review of DMD in its most basic form, while directing the reader to DMDbook and the references therein for a more complete discussion. Assume that we have data snapshots $\{X_{n}\}_{n=1}^{N+1}\subset\mathbb{C}^{M}$ for $N\geq 1$. In the coming applications these snapshots will be vectorized greyscale video frames, and so we make the assumption that they are arranged sequentially. Precisely, each snapshot $X_{n}$ is generated by the same process and sampled at times $t_{n}$ that are evenly spaced, i.e. $h:=t_{n+1}-t_{n}>0$ for all $n=1,\dots,N$. DMD then seeks to identify a matrix $A\in\mathbb{C}^{M\times M}$ so that
@@
Identifying an appropriate DMD matrix is achieved by least squares regression. First, arrange the snapshots into $M\times N$ dimensional data matrices
@@
where $(\lambda_{m},\psi_{m})\in\mathbb{C}\times\mathbb{C}^{M}$ are the $M$ eigenpairs of $A$, and the coefficients $c_{m}$ are given by $\Psi^{-1}X_{1}$, where $\Psi=[\psi_{1}\ \psi_{2}\ \dots\ \psi_{M}]$. Expansion (4) provides a space-time separation of variables for the data with eigenvectors describing macroscopic patterns, or coherent structures, that can only grow, decay, or oscillate in time according to their eigenvalue. Importantly, notice that eigenvectors with eigenvalues $|\lambda_{m}|\approx 1$ are modes that change little in time, giving way to the interpretation that they are the ‘background’ in the data. To better separate our background and foreground modes we sort according to the ‘continuous-time’ eigenvalues, given by
@@ 2.2 Compressed Dynamic Mode Decomposition
In the case where snapshots are high-dimensional ($M\gg 1$), as in the case of high-resolution videos, applying DMD is computationally expensive due to the pseudo-inversion required to identify $A$. This computational bottleneck can be overcome using compressed DMD (cDMD) erichson2019compressed which we incorporate into our method below. Begin by fixing an integer $p\ll M$ and generate a random measurement matrix $C\in\mathbb{C}^{p\times M}$. As explained in erichson2019compressed , this matrix can be constructed componentwise via sampling a uniform or Gaussian distribution, among other choices (we employ the former, which is expressed as $C=\text{rand}(p,M)$ in MATLAB). With this matrix we compress our data matrices (2) into $p\times N$ matrices
@@
To further eliminate redundancies in our data, we can rank-reduce the data using the singular value decomposition tu2014dmd . This is achieved through the reduced SVD $X^{\prime}=U_{r}\Sigma_{r}V^{*}_{r}$, where $U\in\mathbb{C}^{p\times r}$, $\Sigma\in\mathbb{C}^{r\times r},V\in\mathbb{C}^{N\times r}$ with $r\ll N$ and $r\leq p$, and the subscript $r$ denotes only the first $r$ columns and/or rows of each matrix in the SVD of $X^{\prime}$. This leads to an $r\times r$ DMD matrix given by
@@
Information from the compressed and rank-reduced matrix $\tilde{A}$ can be leveraged to obtain information about the original data in $X$ and $Y$. This comes from first letting $Q\in\mathbb{C}^{r\times r}$ be a matrix of eigenvectors of $\tilde{A}$ with the diagonal matrix of eigenvalues $\Lambda$ satisfying $\tilde{A}Q=Q\Lambda$. Full $M$-dimensional DMD eigenvectors are recovered through the $M\times r$ dimensional matrix
@@
where we notably use the original data matrix $Y$. The columns of $\Phi$ contain the DMD modes to construct the space-time expansion (4), with corresponding eigenvalues on the diagonal of $\Lambda$.
Overall, cDMD proves to be much more computationally efficient due a smaller SVD to calculate. The only additional computational costs comes from generating the measurement matrix $C$ and computing the respective matrix products $X^{\prime}$ and $Y^{\prime}$ as in (7). With our method of motion detection for video surveillance, generating $C$ comes as a one-time upfront cost that reduces numerous computations over time. This is because we will apply DMD to incoming windows of video data, each of which is compressed according to the same compression matrix, thus reducing computation within each window.
@@ 2.3 Sliding Window DMD
This idea of applying DMD in real-time to segments of video data is inspired by the sliding window DMD algorithm dylewsky2019dynamic , which was originally devised to better resolve multiscale dynamics. In this framework, snapshots are separated into overlapping, consecutive chunks and DMD is applied on each. To be specific, the $N+1$ snapshots of the video are divided into windows of size $T>0$ so that each window is identified with a value $k\in\{1,\dots,N-T\}$. The $k$th data matrices have the form
@@ 2.4 Background Subtraction with DMD
An impressively successful application of DMD is for performing background separation from video han2022quaternion ; kutz2015multi ; grosek2014dynamic ; krake2022efficient . The method takes advantage of the linearity of DMD by separating out the fast- and slow-moving modes in (4). Recalling that the eigenvectors of the DMD matrix correspond to coherent structures in the data, eigenvalues with modulus near 1 are associated with components that exhibit little change over time. In the context of a video filmed with a fixed camera position, the portion of the video with unchanging dynamics constitutes the background, while the remaining eigenvectors are associated with the foreground. To this end, we are able to approximate the background by considering (4) summed over only eigenvectors with eigenvalues near 1, sometimes referred to as the low-rank component of the data. Summing over the foreground constitutes the sparse component.
@@
To properly separate out low-rank and sparse components from video data $W$ we perform the following. First, define a threshold $\varepsilon>0$ that separates out fast and slow frequencies (5). We set
@@
and note that with real video data it is often the case that the sum defining $\tilde{L}$ is over only a single eigenvector. This comes from a clear distinction between frequencies with $\omega_{k}\approx 0$ and those whose frequencies are bounded away from zero, providing a robustness with respect to the choice of $\varepsilon$. The reader is referred to Figure 3 for a demonstration of this. Following grosek2014dynamic , to maintain that the sparse component is real-valued, we set
@@
where $X_{n}$ is the $n$th video frame and $|\tilde{L}|$ is the element-wise modulus of $L$. However, this process may result in negative pixel values in $\tilde{S}$ and so we define the residual $R$ to be a matrix of all such negative values in $\tilde{S}$. This leads to the true background $L$ and foreground $S$ components of the video
@@
We empirically observed that defining $L=|\tilde{L}|$ resulted in improved background separation compared to defining $L$ as in (14), and so we used this formulation. This method works well, as evidenced by Figure 1. In practice, this procedure is applied in tandem with sliding window DMD on the reconstruction as in (11). In this instance however, we only sum over the slow and fast frequencies in the eigenvalue/eigenvector expansions (6) to reconstruct the background and foreground respectively.

### v2

# Real-Time Motion Detection Using Dynamic Mode Decomposition

@@ Abstract
Dynamic Mode Decomposition (DMD) is a numerical method that seeks to fit time-series data to a linear dynamical system. In doing so, DMD decomposes dynamic data into spatially coherent modes that evolve in time according to exponential growth/decay or with a fixed frequency of oscillation. A prolific application of DMD has been to video, where one interprets the high-dimensional pixel space evolving through time as the video plays. In this work, we propose a simple and interpretable motion detection algorithm for streaming video data rooted in DMD. Our method leverages the fact that there exists a correspondence between the evolution of important video features, such as foreground motion, and the eigenvalues of the matrix which results from applying DMD to segments of video. We apply the method to a database of test videos which emulate security footage under varying realistic conditions. Effectiveness is analyzed using receiver operating characteristic curves, while we use cross-validation to optimize the threshold parameter that identifies movement.
@@
dynamic mode decomposition, motion detection, background subtraction, computer vision
@@ 1 Introduction
Dynamic Mode Decomposition (DMD) is a method for forecasting and extracting complicated spatio-temporal behaviour from high-dimensional datasets. First developed by Schmid [39], DMD seeks to identify a matrix $A$ so that advancing the data forward by one timestep is equivalent to multiplication by this matrix [49, 24]. In this way, DMD poses complex movement over space and through time as a linear discrete dynamical system via the matrix $A$. Since linear systems are entirely understood by their spectrum, the eigenvectors of $A$, called the DMD modes, can be seen to represent macroscopic patterns in the dataset that persist through time, either oscillating, growing, or decaying at a fixed rate determined by the associated eigenvalue. DMD has connections to the Koopman operator [38, 8] and has proven to be a useful tool in identifying coherent structures in fluid mechanics [41, 30, 10] and neuroscience [7, 34, 42], as well as to develop financial trading strategies [29], among numerous other applications; see [24, 40] and the references therein. Improvements on the method include building in known physics [2, 31], accounting for irregular timesteps and noisy measurements [26, 1, 20], and to incorporate the effect of control [35, 16].
The utility of DMD has also made its way into computer vision. This includes employing DMD for edge detection in images [3], salient region detection [45, 43, 44], and performing image classification [36]. Most important to the work herein, DMD has been applied to separate out the background and foreground in videos, allowing for object tracking [22, 17, 25, 14, 19, 18]. Roughly, this comes from the fact that eigenvectors of the DMD matrix $A$ with eigenvalues having modulus near 1 represent patterns in the (vectorized) video data that exhibit little change from frame to frame. Thus, these eigenvectors can be seen as representing the background of the video data, while the remaining eigenvectors fill in the foreground motion. Importantly, projecting the video onto the background/foreground DMD modes constitutes a separation of these distinct components of the video. The result is then two videos: one of a nearly static background and another of just the isolated motion, as exemplified in Figure 1.
@@
Figure 1: Screenshots of DMD being used to separate out background and foreground in video data. The video is taken from the Background Models Challenge database [50].
@@
Figure 2: A visual overview of real-time motion detection and tracking with DMD. Sequential collections of video frames, or windows, are fed in and compressed to reduce computational complexity. Then, DMD is applied to identify spatio-temporal features within the windows. Spikes in the spectrum of the DMD matrices over the windows indicate movement in the video, thus detecting events as they occur in real-time.
@@
Our contribution in this paper is to present a method of real-time motion detection in streaming video data using DMD. To do this we apply DMD not to the entire video all at once, but to short windows as they are fed in. Precisely, we employ sliding window DMD [13, 6] to streaming video data to examine the dominant timescales in each segment of the video. These timescales are revealed by examining the modulus of the logarithm of the eigenvalues for the DMD matrix within each window. Sudden spikes in the DMD spectrum denote fast timescale movement in the foreground, allowing for motion detection. A visual overview of our method can be found in Figure 2, while a proper explanation of the method can be found in Section 3 below. Once movement has been detected, one can then use the DMD matrix to isolate the foreground motion by subtracting out the background. Computation time is accelerated by compressing the frames of the video [14], resulting in a computationally cheap and efficient method for real-time motion detection in streaming video data.
@@
Motion detection is a classic problem in computer vision with many complementary and competing methods. The simplest method identifies temporal differences [9, 51, 23], amounting to searching for significant change in the pixels from one frame to the next. This method is sensitive to abrupt changes in illumination, shadows, and repetitive motions, while also having the drawback that it cannot accurately extract the moving object boundaries. Alternatively, the work [48] transforms the frames into Fourier space where object movement can be shown to create a non-vertical structure in the 2D Fourier transform. While again being an effective method, it suffers from a high computational cost. Computational cost can be brought down using the cosine transform [33], for which graph cuts can be employed to isolate movement in the video. Graph cuts come with their own well-documented limitations (see [5] and the references therein), including significant memory usage. Another drawback of coupling motion detection algorithms with graph cuts is the simple fact that identifying and isolating movement becomes a two-step task. A major advantage of our method herein is that DMD provides the necessary tools to both identify movement (through eigenvalues) and isolate it (through eigenvectors).
@@
The review [46] compares eight methods for real-time video surveillance of varying complexity. It is found that while simpler methods, such as examining temporal differences, are fast, they often fail to handle many of the complexities encountered in real-world video data. As expected, better performing methods come with higher computation time and a more complicated implementation. This includes a recent interest in applying neural networks to motion detection tasks [15, 28, 12, 27]. This however comes at the cost of requiring a lot of training data to initialize the method and tedious hyperparameter tuning to obtain optimal results. This means that reproducing results is often difficult, while significant time and effort can be spent both curating training data and training the network itself. Even though we cannot claim that our DMD-based method herein can outperform state-of-the-art neural network or Gaussian mixture models [4, 37, 32], it offers simplicity and speed in its implementation, while also being backed up by dynamical systems theory to justify its performance. Fundamentally, DMD is a multiresolution analysis method that simultaneously accounts for spatial and temporal coherencies. It identifies macroscopic patterns in the data and is more robust to scale overlap than related methods [13], making it optimally suited for the task of motion detection. Furthermore, our array of test videos demonstrate that our method can be applied to situations that cause difficulties for others, such as multi-object tracking [11] and illumination variation.
@@
The paper is organized as follows. In Section 2 we provide an overview of DMD (Section 2.1, 2.2, 2.3) and its capabilities for background subtraction (Section 2.4), followed by a presentation of our own technique for motion detection in video (Section 3). In Section 4, we apply this method to a dataset of videos (Section 4.1) and discuss its performance (Section 4.2). The performance is then evaluated through the lens of receiver operating characteristic curves (Section 4.3), and a strategy to optimize a key parameter involving $k$-fold cross-validation is presented in Section 4.4. In Section 4.5 we showcase the performance of the method on a publicly available benchmark video database, and provide concluding remarks in Section 5.
@@ 2 Background on Dynamic Mode Decomposition
In this section we provide an overview of all the relevant aspects of DMD for our method. We begin in Section 2.1 with a presentation of DMD in its simplest form. We then turn to a review of DMD with compression in Section 2.2, sliding window DMD in Section 2.3, and background subtraction in Section 2.4.
@@ 2.1 Dynamic Mode Decomposition
We begin with a review of DMD in its most basic form, while directing the reader to [24] and the references therein for a more complete discussion. Assume that we have data snapshots $\{X_{n}\}_{n=1}^{N+1}\subset\mathbb{C}^{M}$ for $N\geq 1$. In the coming applications these snapshots will be vectorized greyscale video frames, and so we make the assumption that they are arranged sequentially. Precisely, each snapshot $X_{n}$ is generated by the same process and sampled at times $t_{n}$ that are evenly spaced, i.e., $h:=t_{n+1}-t_{n}>0$ for all $n=1,\dots,N$. DMD then seeks to identify a matrix $A\in\mathbb{C}^{M\times M}$ so that
@@
Identifying an appropriate DMD matrix is achieved by least squares regression. First, arrange the snapshots into $M\times N$ data matrices
@@
where $(\lambda_{m},\psi_{m})\in\mathbb{C}\times\mathbb{C}^{M}$ are the $M$ eigenpairs of $A$, and the coefficients $c_{m}$ are given by $\Psi^{-1}X_{1}$, where $\Psi=[\psi_{1}\ \psi_{2}\ \dots\ \psi_{M}]$. Expansion (4) provides a space-time separation of variables for the data with eigenvectors describing macroscopic patterns, or coherent structures, that can only grow, decay, or oscillate in time according to their eigenvalue. Importantly, notice that eigenvectors with eigenvalues $|\lambda_{m}|\approx 1$ are modes that change little in time, leading to the interpretation that they are the ‘background’ in the data. To better separate background and foreground modes we sort according to the ‘continuous-time’ eigenvalues, given by
@@ 2.2 Compressed Dynamic Mode Decomposition
When snapshots are high-dimensional ($M\gg 1$), as in the case of high-resolution videos, applying DMD is computationally expensive due to the pseudo-inversion required to identify $A$. This computational bottleneck can be overcome using compressed DMD (cDMD) [14] which we incorporate into our method below. Begin by fixing an integer $p\ll M$ and generate a random measurement matrix $C\in\mathbb{C}^{p\times M}$. As explained in [14], this matrix can be constructed componentwise via sampling a uniform or Gaussian distribution, among other choices (we employ the former, which is expressed as $C=\text{rand}(p,M)$ in MATLAB). With this matrix we compress our data matrices (2) into $p\times N$ matrices
@@
To further eliminate redundancies in our data, we can rank-reduce the data using the Singular Value Decomposition (SVD) [49]. This is achieved through the reduced SVD $X^{\prime}=U_{r}\Sigma_{r}V^{*}_{r}$, where $U\in\mathbb{C}^{p\times r}$, $\Sigma\in\mathbb{C}^{r\times r},V\in\mathbb{C}^{N\times r}$, with $r\ll N$ and $r\leq p$, and the subscript $r$ denotes only the first $r$ columns and/or rows of each matrix in the SVD of $X^{\prime}$. This leads to an $r\times r$ DMD matrix given by
@@
Information from the compressed and rank-reduced matrix $\tilde{A}$ can be leveraged to obtain information about the original data in $X$ and $Y$. This comes from first letting $Q\in\mathbb{C}^{r\times r}$ be a matrix of eigenvectors of $\tilde{A}$ with the diagonal matrix of eigenvalues $\Lambda$ satisfying $\tilde{A}Q=Q\Lambda$. Full $M$-dimensional DMD eigenvectors are then recovered through the $M\times r$ matrix
@@
where we notably use the original data matrix $Y$. The columns of $\Phi$ contain the DMD modes needed to construct the space-time expansion (4), with corresponding eigenvalues on the diagonal of $\Lambda$.
Overall, cDMD proves to be much more computationally efficient due a smaller SVD to calculate. The only additional computational costs comes from generating the measurement matrix $C$ and computing the respective matrix products $X^{\prime}$ and $Y^{\prime}$ as in (7). With our method of motion detection for video surveillance, generating $C$ comes as a one-time upfront cost that reduces numerous computations over time. This is because we will apply DMD to incoming windows of video data, each of which is compressed using the same compression matrix, thus reducing computation within each window.
@@ 2.3 Sliding Window DMD
This idea of applying DMD in real-time to segments of video data is inspired by the sliding window DMD algorithm [13], which was originally devised to better resolve multiscale dynamics. In this framework, snapshots are separated into overlapping, consecutive chunks and DMD is applied on each. To be specific, the $N+1$ snapshots of the video are divided into windows of size $T>0$ so that each window is identified with a value $k\in\{1,\dots,N-T\}$. The $k$th data matrices have the form
@@ 2.4 Background Subtraction with DMD
An impressively successful application of DMD is background separation in video data [18, 25, 17, 22]. The method takes advantage of the linearity of DMD by separating out the fast- and slow-moving modes in (4). Recalling that the eigenvectors of the DMD matrix correspond to coherent structures in the data, eigenvalues with modulus near 1 are associated with components that exhibit little change over time. In the context of a video filmed with a fixed camera position, the portion of the video with unchanging dynamics constitutes the background, while the remaining eigenvectors are associated with the foreground. To this end, we are able to approximate the background by considering (4) summed over only eigenvectors with eigenvalues near 1, sometimes referred to as the low-rank component of the data. Summing over the foreground constitutes the sparse component.
@@
To properly separate out low-rank and sparse components from video data we perform the following. First, define a threshold $\varepsilon>0$ that separates out fast and slow frequencies (5). We set
@@
and note that with real video data it is often the case that the sum defining $\tilde{L}$ is over only a single eigenvector. This comes from a clear distinction between frequencies with $\omega_{k}\approx 0$ and those whose frequencies are bounded away from zero, providing robustness with respect to the choice of $\varepsilon$. The reader is referred to Figure 3 for a demonstration of this. Following [17], to ensure that the sparse component is real-valued, we set
@@
where $X_{n}$ is the $n$th video frame and $|\tilde{L}|$ is the element-wise modulus of $L$. However, this process may result in negative pixel values in $\tilde{S}$ and so we define the residual $R$ to be a matrix of all such negative values in $\tilde{S}$. This leads to the true background $L$ and foreground $S$ components of the video, defined as
@@
We empirically observed that defining $L=|\tilde{L}|$ resulted in improved background separation compared to defining $L$ as in (14), and so we used this formulation. This method works well, as evidenced by Figure 1. In practice, this procedure is applied in tandem with sliding window DMD on the reconstruction as in (11). In this instance however, we only sum over the slow and fast frequencies in the eigenvalue/eigenvector expansions (6) to reconstruct the background and foreground, respectively.